OpenStudy (anonymous):

Write an equation of a line in point-slope form that passes through the point (1,2) and whose slope is 3.

jimthompson5910 (jim_thompson5910):

\[\large y-y_{1}=m(x-x_{1})\] \[\large y-2=3(x-1)\] And that's as far as you go because the last equation above is in point-slope form.